Six artificial satellites circle a space station a


Six artificial satellites circle a space station a



Please rank in order from largest to smallest ie:


3rd> 1st> 5th


For full points.
Answer #1

a = v^2/R

a1 = 40^2/10000 = 0.16 m/s^2

a2 = 80^2/2500 = 2.56 m/s^2

a3 = 120^2/5000 = 2.88 m/s^2

a4 = 160^2/2500 = 10.24 m/s^2

a5 = 160^2/5000 = 5.12 m/s^2

a6 = 80^2/10000 = 0.64 m/s^2


clearly,

a4 > a5 >a3 > a2 > a6 > a1

so,

4th > 5th > 3rd > 2nd > 6th > 1st
Answer #2

1) For uniform circular motion, Centripetal Force is given by either of two equations:

F = mv^2/r
F = w^2rm

m = mass
r = radius
w = "omega" = angular velocity
v = tangential velocity

All of these satellites have the same angular velocity, so you just need to order them by decreasing mL to get the list in order from highest to lowest.

A = 8000000

B = 1000000

C = 1000000

D = 250000

E = 1000000

F = 3000000

4th > 5th > 3rd > 2nd > 6th > 1st
Answer #3

4 > 5 > 3 > 2 >6>1